OpenStudy (anonymous):

How do you find the value of x that makes the statement true. 3/7x+3-x+3=2

OpenStudy (anonymous):

Can you use parentheses to make your question clearer?

OpenStudy (muzzack):

One solution was found : \(\color{blue}{\huge\cal\ x = 7~}\)

OpenStudy (anonymous):

(3/7x+3)-(x+3)=2 better?

OpenStudy (zzr0ck3r):

\(\frac{3}{7x+3}-(x+3)=2?\)

OpenStudy (anonymous):

\[(\frac{ 3 }{ 7x }+3)-(x+3)=2\]

OpenStudy (zzr0ck3r):

how you wrote it does not help.... (3/7x+3)-(x+3)=2 is the same as \(\frac{3}{7x}+3-x-3\)

OpenStudy (zzr0ck3r):

maybe draw it....